OpenStudy (anonymous):

Explain how to find the sum positive five plus negative nineusing chips, a number line, or the rules for adding integers. Give the answer with your explanation.

OpenStudy (anonymous):

@chosenmatt

OpenStudy (anonymous):

@leonardo0430

OpenStudy (anonymous):

@jordanloveangel

OpenStudy (jordanloveangel):

.Explain how to find the sum positive three plus negative nine using chips, a number line, or the rules for adding integers. Give the answer with your explanation. let's use the number line! Start at zero, go to the right 3 units (positive three), then go to the left 9 units (negative nine). Stop and look at the number line, you are at -6 position on the number line. That means the sum of +3 and -9 is -6

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

??????????????????????????????????????????????????????????????

OpenStudy (jordanloveangel):

did this help @brandonsgirl

OpenStudy (anonymous):

sorta

OpenStudy (jordanloveangel):

5+(−9)=5−9=?

OpenStudy (anonymous):

Write a two-step equation that needs the steps of subtraction then multiplication in order to be solved, and then solve it. Show your work.

OpenStudy (anonymous):

@jordanloveangel

OpenStudy (jordanloveangel):

1 sec

OpenStudy (jordanloveangel):

a 2 step equation means that it must involve 2 steps. So, you could have an equation like (x-y)*(z). The x,y, and z stand for numbers you would input. To solve, you need to do the subtraction (x-y) first then multiply it by the "z" value. Use order of operations (PEMDAS) to solve out longer equations.

OpenStudy (anonymous):

thank you XD XD XD

OpenStudy (jordanloveangel):

any time u got more

OpenStudy (anonymous):

i think

OpenStudy (jordanloveangel):

oh kk if u do just PM me or tag me

OpenStudy (anonymous):

no more and i will